9) State the Divergence Theorem and verify it for me re, y, z) =< 3x, xy, 2xz> and the unit cube F defined by verify it fo
Answer #1

Statement of Gauss Divergence theorem Let S be the closed surface enclosing a volume v. It Ē is a continuously . differentiabI dive du = 2 jtesia divp dv : 1-0) dz z=0 (1-0) . div F dr = 2 O YA > verification we will cal erlate the value of Pfinds ov=3 tt-o] JS er ds as tas-- - for 52 - OCRB unit outward drawn normal ñ=-1 =0, ds = dydz : 0242 047 41 F.ñ = 3% -0 250 d o dydFor s4 - OASC 2-O -K s - O4*10cy1 fin = -272 EN=0 Z0 En ds = 1 To dondy =0 u sasa x = y= D Ends = 0 - (V) ds a da dz Foy 5 =For SG= OBQA, Y=0, n=-j ds = dx dz Ocnel, Oczel En = -xy Fń - y=0 Finds = o dadz = 0 X=0 220 s Ende =0 @ Foods = $58 dat seti